Judy wants to invest $50,000. Part of it is invested in a fund that pays 12.5% and the remainder in one that pays 14%. Find the amount invested at each rate if the annual income from the two is $6,640.
Answer by mananth(16946)

Investment I 12.50% per annum
Investment II 14.00% per annum

x + y 50000 ------------------------1
12.50% x + 14.00% y = $6,640.00
Multiply by 100
12.5 x + 14 y = $664,000.00 --------2

Multiply (1) by -12.5
we get

-12.5 x -12.5 y = -625000.00

Add this to (2)

0 x 1.5 y = $39,000.00

divide by 1.5

y= $26,000.00 investment at 14.00%
Balance $24,000.00 investment at 12.50%
CHECK
$24,000.00------- $3,000.00
$26,000.00------- $3,640.00
Total $6,640.00
m.ananth@hotmail.ca